OpenStudy (anonymous):

"A merchant sells three different sizes of canned tomatoes. A large can costs as much as 5 medium cans, or 7 small cans. If a customer buys an equl number of small and large cans of tomatoes for the exact amount of money that would buy 200 medium cans, how many small cans will she buy? " HELP! HOW DO I SET THIS EQUATION UP?

OpenStudy (anonymous):

@ MathsoRusty are u familiar with linear equations and solving simultaneous equations?

OpenStudy (cwrw238):

let L = cost of 1 large, M = cost 1 medium and S = cost of 1 small can now we can write some equations relating these variables: L = 5M L = 7S 200M = xL + xS where x = number of small and large cans

OpenStudy (anonymous):

From cwrw238 argument: let 200M=xL+xS...........(1) L=5M.....(2) L=7S.....(3) (2)=(3) therefore 5M=7S from this S=5M/7.......(4) let substitute (2)and (4) into (1) 200M=x(5M)+x(5M/7) from this you'll get x=35

OpenStudy (anonymous):

is that cool so far?